OpenStudy (anonymous):

i know the answer A model for the data is given by the following. t=38m+16965/10(m+5000) Use the model to approximate the mass of an object when the average time for one oscillation is 1.058 seconds. Round your answer to the nearest gram.

OpenStudy (anonymous):

the answer is 1311grams

OpenStudy (amistre64):

1.058 = (38m+16965)/10(m+5000) ; and solve for m

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

1.058 = (38m+16965)/10(m+5000) / *(m+5000) 1.058m + 1.058(5000) = 38m +16965 ; switch values to combine like terms 5290 - 16965 = 36.942m ; /36.942 -316.03... = m well its work if I could decipher the equation you typed

OpenStudy (amistre64):

i dropped a 10 lol

OpenStudy (anonymous):

don't worry about the one, i got another problem similar to this one, but dont got the answer,

OpenStudy (anonymous):

A model for the data is given by the following. Use the model to approximate the mass of an object when the average time for one oscillation is 1.031 seconds. Round your answer to the nearest gram. t=35m+16965/10(m+5000)

OpenStudy (amistre64):

1.058 = (38m+16965)/10(m+5000) / *(10m+50000) 10.58m + 1.058(50000) = 38m +16965 52900 - 16965 = 27.42m 1310.53... = m

OpenStudy (anonymous):

yeah, got the with the calculator

OpenStudy (anonymous):

should it not be 10m? in the first step? after you distribute the 10?

OpenStudy (amistre64):

1400.77... is what I get with the new one

OpenStudy (amistre64):

teh old one: 1.058(10m) = 10.58m

OpenStudy (anonymous):

no, you had 10(m+5000)